### v2

# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

Abstract
Spatial reasoning is a fundamental aspect of human intelligence. One key concept in spatial cognition is the Frame of Reference (FoR), which identifies the perspective of spatial expressions. Despite its significance, FoR has received limited attention in AI models that need spatial intelligence. There is a lack of dedicated benchmarks and in-depth evaluation of large language models (LLMs) in this area. To address this issue, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark, designed to assess FoR comprehension in LLMs. We evaluate LLMs on answering questions that require FoR comprehension and layout generation in text-to-image models using FoREST. Our results reveal a notable performance gap across different FoR classes in various LLMs, affecting their ability to generate accurate layouts for text-to-image generation. This highlights critical shortcomings in FoR comprehension. To improve FoR understanding, we propose Spatial-Guided prompting, which improves LLMs’ ability to extract essential spatial concepts. Our proposed method improves overall performance across spatial reasoning tasks.
FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks
Tanawan Premsri Department of Computer Science Michigan State University premsrit@msu.edu                        Parisa Kordjamshidi Department of Computer Science Michigan State University kordjams@msu.edu
1 Introduction
Spatial reasoning plays a significant role in human cognition and daily activities. It is also a crucial aspect in many AI problems, including language grounding (Zhang and Kordjamshidi, 2022; Yang et al., 2024), navigation (Yamada et al., 2024), computer vision (Liu et al., 2023; Chen et al., 2024), medical domain (Gong et al., 2024), and image generation (Gokhale et al., 2023). One key concept in spatial reasoning is the Frame of Reference (FoR), which identifies the perspective of spatial expressions. FoR has been studied extensively in cognitive linguistics (Edmonds-Wathen, 2012; Vukovic and Williams, 2015). Levinson (2003) initially defines three FoR classes: relative, based on the observer’s perspective; intrinsic, based on an inherent feature of the reference object; and absolute, using environmental cues like cardinal directions -See Figure 1. This framework was expanded by  Tenbrink (2011) to create a more comprehensive framework, serving as the basis of our work. Understanding FoR is important for many applications, especially in embodied AI. In such applications, an agent must simultaneously comprehend multiple perspectives, including the one from the instruction giver and from the instruction follower, to communicate and perform tasks effectively. However, recent spatial evaluation benchmarks have largely overlooked FoR. For example, the text-based benchmarks Shi et al. (2022); Mirzaee and Kordjamshidi (2022); Rizvi et al. (2024) and text-to-images benchmarks (Gokhale et al., 2023; Huang et al., 2023; Cho et al., 2023a, b) assume a fixed perspective for all spatial expressions. This inherent bias limits situated spatial reasoning, restricting adaptability in interactive environments where perspectives can change.
Figure 1: Illustration of FoR classes. The cat is the locatum, the car is the relatum, and the arrow denotes the perspective.
To systematically investigate the role of FoR in spatial understanding and create a new resource, that is, Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST), FoREST is designed to evaluate models’ ability to comprehend FoR from textual descriptions and extend this evaluation to grounding and visualization. Our benchmark includes spatial expressions with FoR ambiguity—where multiple FoRs may apply to the described situation—and spatial expressions with only a single valid FoR. This design allows evaluation of the models’ understanding of FoR in both scenarios. We evaluate several LLMs in a QA setting that require FoR understanding and apply the FoR concept in text-to-image models. Our findings reveal performance differences across FoR classes and show that LLMs exhibit bias toward specific FoRs when handling ambiguous cases. This bias extends to layout-diffusion models, which rely on LLM-generated layouts in the image generation pipeline. To enhance FoR comprehension in LLMs, we propose Spatial-Guided prompting, which enables models to analyze and extract additional spatial information, including directional, topological, and distance relations. We demonstrate that incorporating spatial information improves question-answering and layout generation, ultimately enhancing text-to-image generation performance.
Our contribution111code and dataset available at GitHub repository. are summarized as follows, 1. We introduce the FoREST benchmark to evaluate LLMs’ FoR comprehension in a QA setting systematically. 2. We analyze the impact of FoR information on text-to-image generation using multiple diffusion models. 3. We propose a prompting approach that generates spatial information, which can be incorporated into QA and layout diffusion to enhance performance.
2 Spatial Primitives
We review three semantic aspects of spatial information expressed in language: Spatial Roles, Spatial Relations, and Frame of Reference.
Spatial Roles. We focus on two main spatial roles (Kordjamshidi et al., 2010) of Locatum, and Relatum. The locatum is the object described in the spatial expression, while the relatum is the other object used to describe the position of the locatum. An example is a cat is to the left of a dog, where the cat is the locatum, and the dog is the relatum.
Spatial Relations. When dealing with spatial knowledge representation and reasoning, three main relations are often considered, directional, topological, and distance (Hernández, 1994; Cohn and Renz, 2008; Kordjamshidi et al., 2011). Directional describes an object’s direction based on specific coordinates, e.g., left and right. Topological describes the containment between two objects, such as inside. Distance describes qualitative and quantitative relations between entities. Examples of qualitative are far, and quantitative are 3km.
Spatial Frame of Reference. We use four frames of references investigated in the cognitive linguistic studies Tenbrink (2011). These are defined based on the concept of Perspective, which is the origin of a coordinate system to determine the direction. The four frames of reference are defined as follows.
1. External Intrinsic describes a spatial relation from the relatum’s perspective, where the relatum does not contain the locatum. The top-right image in Figure 1 illustrates this with the sentence, A cat is to the right of the car from the car’s perspective.
2. External Relative describes a spatial relation from the observer’s perspective. The top-left image in Figure 1 shows an example with the sentence, A cat is to the left of a car from my perspective.
3. Internal Intrinsic describes a spatial relation from the relatum’s perspective, where the relatum contains the locatum. The bottom-right image in Figure 1 shows this with the sentence, A cat is inside and back of the car from the car’s perspective.
4. Internal Relative describes a spatial relation from the observer’s perspective where the locatum is inside the relatum. The bottom-left image in Figure 1 shows this FoR with the sentence, A cat is inside and to the left of the car from my perspective.
Figure 2: Pipeline for dataset creation, starting from selecting a locatum and relatum from available objects and then applying a spatial template to generate the spatial expression ($T$). FoRs are assigned based on relatum properties. $T$ is categorized based on the number of FoRs. For example, A cat is to the right of a dog (with two possible FoRs: external intrinsic and external relative) belongs to the A-split. Then, its disambiguated version (A cat is to the right of a dog from the dog’s perspective) is added to the C-split. Next, if applicable, a relatum’s orientation is included for visualization and question generation. Finally, Unity3D generates scene configurations, and question-answer pairs are derived from $T$.
3 FoREST Dataset Construction
To systematically evaluate LLM on the frame of reference (FoR) recognition, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark. Each instance in FoREST consists of a spatial context ($T$), a set of corresponding FoR ($FoR$) which is a subset of {external relative, external intrinsic, internal intrinsic, internal relative}, a set of questions and answers ($\{Q,A\}$), and a set of visualizations ($\{I\}$). An example of $T$ is A cat is to the right of a dog. A dog is facing toward the camera. The FoR of $T$ is {external intrinsic, external relative}. A question-answer is $Q$ = Based on the camera’s perspective, where is the cat from the dog’s position?, $A$ = {left, right}. There is an ambiguity in the FoR for this expression. Thus, the answer will be left if the model assumes the external relative. Conversely, it will be right if the model assumes the external intrinsic. The visualization of this example is in Figure 2.
3.1 Context Generation
We select two distinct objects—a relatum ($R$) and a locatum ($L$)—from a set of 20 objects and apply them to a Spatial Relation template, <$L$> <spatial relation> <$R$> to generate the context $T$. FoRs for $T$ are determined based on the properties of the selected objects. Depending on the number of possible FoRs, $T$ is categorized as ambiguous (A-split), where multiple FoRs apply, or clear (C-split), where only one FoR is valid. We further augment the C-split with disambiguated spatial expressions derived from the A-split, as shown in Figure 2.
3.2 Categories based on Relatum Properties
Using the FoR classes in Section 2, we identify two key properties contributing to FoR ambiguity. The first property is the relatum’s intrinsic direction. It creates ambiguity between intrinsic and relative FoRs as spatial relations may originate from either relatum’s or observer’s perspectives. The second is the relatum’s affordance as a container. It introduces the ambiguity between internal and external FoR, as spatial relations may refer to either inside or outside of the relatum. Note that containment is defined as the ability of the relatum to contain the locatum, considering both objects’ sizes. Based on these properties, we define four distinct cases: Cow, Box, Car, and Pen.
Case 1: Cow Case. In this case, the selected relatum has intrinsic directions but cannot be the container for the locatum. An example is a cow. In such cases, the relatum potentially provides a perspective for spatial relations. The applicable FoR classes are $FoR$ = {external intrinsic, external relative}. We augment the C-split with expressions of this case but include the perspective to resolve their ambiguity. To specify the perspective, we use predefined templates for augmenting clauses, such as from {relatum}’s perspective for external intrinsic or from the camera’s perspective for external relative. For example, if the context is A cat is to the right of the cow, in the A-split. The counterparts included in the C-split are A cat is to the right of the cow from cow’s perspective. for external intrinsic and A cat is to the right of the cow from my perspective. for external intrinsic.
Case 2: Box Case. The relatum in this case has the property of being a container but lacks intrinsic directions, making the internal FoR applicable. An example is a box. The applicable FoR classes are $FoR$ = {external relative, internal relative}. To include their unambiguous counterparts in the C-split, we specify the topological relation by adding inside for internal relative and outside for external relative. For example, for the sentence A cat is to the right of the box., the unambiguous $T$ with internal relative FoR is A cat is inside and to the right of the box. The counterpart for external relative is A cat is outside and to the right of the box.
Case 3: Car Case. A relatum with an intrinsic direction and container affordance falls into this case, allowing all FoR classes. An obvious example is a car that can be a container with intrinsic directions. The applicable FoR classes are $FoR$ = { external relative, external intrinsic, internal intrinsic, internal relative}. To augment C-split with this case’s disambiguated counterparts, we add perspective and topology information similar to the Cow and Box cases. An example expression for this case is A person is in front of the car. The four disambiguated counterparts to include in the C-split are A person is outside and in front of the car from the car itself. for external intrinsic, A person is outside and in front of the car from the observer. for external relative, A person is inside and in front of the car from the car itself. for internal intrinsic, and A person is inside and in front of the car from the observer. for internal relative.
Case 4: Pen Case. In this case, the relatum lacks both the intrinsic direction and the affordance as a container. An obvious example is a pen. Lacking these two properties, the created context has only one applicable FoR, $FoR$ = {external relative}. Therefore, we can categorize this case into both splits without any modification. An example of such a context is The book is to the left of a pen.
3.3 Context Visualization
In our visualization, complexity arises when the relatum has an intrinsic direction within the intrinsic FoR, as its orientation can complicate the spatial representation. For example, for visualization of A cat is to the right of a dog from the dog’s view, the cat can be placed in different coordinates based on the dog’s orientation. To address this issue, we add a template sentence for each direction, such as <relatum> is facing toward the camera, to specify the relatum’s orientation of all applicable $T$ for visualization and QA. For instance, A cat is to the left of a dog. becomes A cat is to the left of a dog. The dog is facing toward the camera. To avoid occlusion issues, we generate visualizations only for external FoRs, as one object may become invisible in internal FoR classes. We use only expressions in C-split since those have a unique FoR interpretation for visualization. We then create a scene configuration by applying a predefined template, as illustrated in Figure 2. Images are generated using the Unity 3D simulator Juliani et al. (2020), producing four variations per expression $T$ with different backgrounds and object positions. Further details on the simulation process are in Appendix A.2.
3.4 Question-Answering Generation
We generate questions for all generated spatial expressions ($T$). Note that we include the relatum orientation for cases where the relatum has an intrinsic direction, as mentioned in the visualization. Our benchmark includes two types of questions. The first type asks for the spatial relation between two given objects from the camera’s perspective, following predefined templates such as, Based on the camera’s perspective, where is the locatum relative to the relatum’s position? Template variations are made based on GPT4o; details are provided in Appendix A.2.4. The second type of question queries the spatial relation from the relatum’s perspective. This question type follows the same templates but replaces the camera with the relatum. The first type of question is generated for all $T$, while the second type is only generated for $T$ where the relatum has intrinsic direction and a perspective can be defined accordingly. Answers are determined based on the corresponding FoRs, the spatial relation in $T$, and the relatum’s orientation when applicable.
4 Models and Tasks
This paper focuses on Question-Answering and Text-to-Image using the FoREST benchmark to evaluate FoR in spatial reasoning comprehensively. FoREST also supports other tasks, such as FoR identification, detailed in Appendix H.
4.1 Question-Answering (QA)
Task. This QA task evaluates LLMs’ ability to adapt contextual perspectives across different FoRs. The input includes a spatial expression $T$, relatum orientation (if available), and a question $Q$ querying the spatial relation from either the observer’s or the relatum’s perspective. The output is a spatial relation $S$, restricted to {left, right, front, back}. We use the context from both splits for this task.
Zero-shot baseline. We call the LLM with instructions, a spatial context, and a question expecting a spatial relation as the response. The prompt instructs the model to answer the question with one of the candidate spatial relations without explanation.
Few-shot baseline. We create four spatial expressions, each assigned to a single FoR class to prevent bias. Following the steps in Section 3.4, we generate a corresponding question and answer for each. These serve as examples in our few-shot prompting. The input to the model is the instruction, example, spatial context, and the question.
Chain-of-Thought (CoT) baseline (Wei et al., 2023). To create CoT examples, we modify the prompt to require reasoning before answering. We manually craft reasoning explanations with the necessary information for the few-shot examples. The input to the model is the instruction, CoT example, spatial context, and the question.
Human baseline. We conducted a small scope human study with three participants, paid as research assistants, to establish a human baseline for A-split. Each participant is shown 25 randomly sampled QA examples per relatum case, totaling 150 examples. Each includes a textual scene description and corresponding question. See Appendix I for details. We report average accuracy across participants to reflect the human baseline.
Camera perspective Relatum perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
RIAcc. RIAcc. Acc. Acc. Acc. RIAcc. RIAcc. Acc.
Llama3-70B (1) $48.1$ $\mathbf{51.5}$ $62.5$ $\mathbf{58.0}$ $41.6$ $65.5$ $73.3$ $72.5$ $64.3$ $\mathbf{61.0}$ $38.7$ $62.1$ $\mathbf{51.8}$ $47.9$ $61.8$ $62.1$
Llama3-70B (2) $49.1$ $\mathbf{50.5}$ $62.2$ $\mathbf{52.2}$ $47.4$ $64.7$ $85.8$ $85.5$ $65.8$ $\mathbf{59.6}$ $40.1$ $57.1$ $\mathbf{55.5}$ $44.2$ $61.8$ $57.7$
Llama3-70B (3) $49.4$ $\mathbf{50.3}$ $80.7$ $49.4$ $\mathbf{50.3}$ $79.6$ $95.8$ $94.9$ $82.6$ $\mathbf{60.8}$ $39.0$ $77.2$ $\mathbf{55.1}$ $44.6$ $80.9$ $77.7$
Llama3-70B (4) $\mathbf{59.4}$ $40.2$ $73.6$ $\mathbf{57.9}$ $41.7$ $74.8$ $100.0$ $100.0$ $77.5$ $\mathbf{60.6}$ $39.1$ $65.7$ $\mathbf{56.0}$ $43.7$ $67.7$ $66.0$
Qwen2-72B (1) $\mathbf{96.6}$ $2.9$ $95.6$ $\mathbf{95.9}$ $3.6$ $95.0$ $100.0$ $100.0$ $96.1$ $8.8$ $\mathbf{90.6}$ $79.3$ $7.8$ $\mathbf{91.7}$ $83.6$ $79.9$
Qwen2-72B (2) $\mathbf{89.0}$ $10.5$ $84.4$ $\mathbf{85.6}$ $13.9$ $85.5$ $100.0$ $100.0$ $86.8$ $17.7$ $\mathbf{81.8}$ $78.3$ $10.4$ $\mathbf{89.1}$ $86.3$ $79.4$
Qwen2-72B (3) $\mathbf{67.2}$ $32.4$ $88.6$ $\mathbf{62.0}$ $37.6$ $83.4$ $100.0$ $100.0$ $89.6$ $21.3$ $\mathbf{78.3}$ $85.5$ $22.7$ $\mathbf{76.9}$ $83.6$ $85.2$
Qwen2-72B (4) $\mathbf{93.0}$ $6.5$ $90.1$ $\mathbf{94.6}$ $4.9$ $93.3$ $100.0$ $98.6$ $91.7$ $8.2$ $\mathbf{91.2}$ $86.0$ $10.5$ $\mathbf{89.0}$ $87.4$ $86.2$
Qwen2VL-72B (1) $49.5$ $\mathbf{50.5}$ $78.3$ $\mathbf{51.6}$ $48.4$ $80.0$ $98.3$ $96.9$ $81.2$ $41.0$ $\mathbf{59.0}$ $55.4$ $44.7$ $\mathbf{55.3}$ $59.0$ $56.0$
Qwen2VL-72B (2) $40.9$ $\mathbf{59.1}$ $89.4$ $44.7$ $\mathbf{55.3}$ $79.0$ $100.0$ $100.0$ $89.6$ $32.7$ $\mathbf{67.3}$ $66.9$ $28.5$ $\mathbf{71.5}$ $67.0$ $66.9$
Qwen2VL-72B (3) $\mathbf{63.5}$ $36.5$ $84.0$ $\mathbf{72.2}$ $27.8$ $84.5$ $100.0$ $100.0$ $86.3$ $\mathbf{51.0}$ $49.0$ $77.7$ $\mathbf{54.6}$ $45.4$ $82.8$ $78.4$
Qwen2VL-72B (4) $\mathbf{50.5}$ $49.5$ $78.3$ $\mathbf{59.5}$ $40.5$ $67.4$ $98.3$ $99.6$ $79.9$ $23.5$ $\mathbf{76.5}$ $78.2$ $33.0$ $\mathbf{67.0}$ $67.7$ $76.7$
GPT-4o (1) $\mathbf{84.3}$ $15.3$ $94.5$ $\mathbf{88.5}$ $11.0$ $97.3$ $99.2$ $99.8$ $95.6$ $21.6$ $\mathbf{78.0}$ $91.6$ $16.1$ $\mathbf{83.5}$ $90.5$ $91.4$
GPT-4o (2) $\mathbf{69.0}$ $30.6$ $76.6$ $\mathbf{80.3}$ $19.2$ $89.5$ $100.0$ $100.0$ $81.5$ $29.0$ $\mathbf{70.5}$ $74.7$ $30.9$ $\mathbf{68.7}$ $77.5$ $75.1$
GPT-4o (3) $41.5$ $\mathbf{58.3}$ $92.3$ $38.2$ $\mathbf{61.6}$ $91.0$ $100.0$ $99.8$ $93.2$ $33.9$ $\mathbf{65.8}$ $93.9$ $32.0$ $\mathbf{67.6}$ $93.9$ $93.9$
GPT-4o (4) $26.0$ $\mathbf{73.9}$ $79.2$ $27.7$ $\mathbf{72.1}$ $79.4$ $96.7$ $94.3$ $81.4$ $16.2$ $\mathbf{83.4}$ $95.5$ $19.2$ $\mathbf{80.4}$ $94.8$ $95.4$
Human-baseline $36.6$ $\mathbf{63.4}$ $90.7$ $27.8$ $\mathbf{72.2}$ $96.0$ $72.0$ $82.7$ $85.3$ $41.4$ $\mathbf{58.6}$ $97.3$ $36.1$ $\mathbf{63.9}$ $96.0$ $96.7$
Table 1: QA accuracy in the A-Split. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
4.2 Text-To-Image (T2I)
Task. This task assesses the diffusion models’ ability to consider FoR by evaluating their generated images. The input is a spatial expression, $T$, and the output is a generated image ($I$). C and A splits with external FoRs are used for this task.
Stable Diffusion Models. We use the stable diffusion models as the baseline for the T2I task. This model only needs the scene description as input.
Layout Diffusion Models. This model operates in two phases: text-to-layout and layout-to-image. Given that LLMs can generate the bounding box layout (Cho et al., 2023b), we provide them with instructions and $T$ to create the layout. The layout consists of bounding box coordinates for each object in the format of {object: $[x,y,w,h]$}, where $x$ and $y$ denote the starting point and $h$ and $w$ denote the height and width. The bounding box coordinates and $T$ are then passed to the layout-to-image model to produce the final image, $I$.
4.3 Spatial-Guide (SG) Prompting
We hypothesize that the spatial relation types and FoR classes defined in Section 2 can improve QA and layout generation. For example, the external intrinsic FoR emphasizes that spatial relations originate from the relatum’s perspective. To leverage this, we propose SG prompting, an additional step applied before QA or layout generation. This step extracts spatial information, including direction, topology, distance, and the FoR, from the spatial expression $T$, as supplementary input to guide LLMs in QA or layout generation. We manually craft four examples covering these aspects. First, we specify the perspective for directional relations, e.g., left relative to the observer, to distinguish intrinsic from relative FoR. Next, we indicate whether the locatum is inside or outside the relatum for topological relations to differentiate internal from external FoR. Lastly, we provide an estimated quantitative distance to support topological and directional relation identification, e.g., far. These examples are provided in a few-shot setting to guide the model in automatically extracting such information.
5 Experimental Results
Camera perspective Relatum perspective
Model ER (CP) EI (RP) II (RP) IR (CP) Avg. ER (CP) EI (RP) II (RP) IR (CP) Avg.
Llama3-70B (0-shot) $44.8$ $38.4$ $39.7$ $54.4$ $42.6$ $42.2$ $47.1$ $62.5$ $34.4$ $45.1$
Llama3-70B (4-shot) $43.0$ $40.0$ $39.1$ $47.3$ $41.9$ $41.8$ $60.9$ $77.7$ $35.2$ $52.0$
Llama3-70B (CoT) $57.8$ $46.1$ $44.7$ $46.0$ $51.5$ $\mathbf{55.5}$ $56.8$ $71.5$ $49.0$ $56.6$
Llama3-70B (SG+CoT) $47.6$ $42.9$ $50.0$ $35.6$ $45.0$ $55.4$ $64.5$ $75.0$ $47.1$ $60.1$
Qwen2-72B (0-shot) $94.5$ $35.2$ $31.8$ $93.2$ $66.9$ $28.7$ $89.3$ $93.6$ $23.8$ $59.0$
Qwen2-72B (4-shot) $90.2$ $39.5$ $39.1$ $68.5$ $65.3$ $33.5$ $92.1$ $94.0$ $29.5$ $62.7$
Qwen2-72B (CoT) $81.4$ $57.4$ $58.6$ $62.5$ $69.1$ $39.5$ $83.7$ $85.2$ $37.7$ $61.6$
Qwen2-72B (SG+CoT) $97.6$ $42.5$ $31.3$ $\mathbf{93.8}$ $71.4$ $42.8$ $86.6$ $92.0$ $34.0$ $64.5$
Qwen2VL-72B (0-shot) $68.3$ $40.2$ $42.4$ $75.3$ $56.0$ $37.4$ $60.1$ $82.2$ $35.5$ $50.0$
Qwen2VL-72B (4-shot) $78.4$ $41.0$ $44.5$ $66.1$ $60.5$ $38.0$ $85.4$ $91.2$ $29.9$ $61.6$
Qwen2VL-72B (CoT) $58.4$ $62.4$ $68.4$ $40.7$ $59.3$ $48.6$ $43.5$ $36.1$ $\mathbf{61.9}$ $46.4$
Qwen2VL-72B (SG+CoT) $\mathbf{99.0}$ $50.0$ $47.5$ $93.0$ $75.9$ $27.3$ $37.3$ $37.1$ $34.0$ $32.7$
GPT-4o (0-shot) $79.7$ $45.1$ $39.5$ $90.2$ $64.2$ $46.9$ $88.5$ $98.2$ $34.8$ $67.5$
GPT-4o (4-shot) $68.0$ $52.6$ $60.7$ $74.1$ $61.8$ $44.9$ $\mathbf{98.2}$ $\mathbf{100.0}$ $37.5$ $71.2$
GPT-4o (CoT) $81.7$ $\mathbf{76.1}$ $\mathbf{82.4}$ $71.5$ $78.8$ $53.0$ $91.1$ $90.6$ $50.8$ $\mathbf{71.9}$
GPT-4o (SG+CoT) $97.9$ $72.2$ $72.7$ $93.4$ $\mathbf{85.8}$ $48.9$ $96.3$ $95.9$ $36.1$ $71.8$
Table 2: QA accuracy in the C-Split across various LLMs. ER, EI, II, and IR denote external relative, external intrinsic, internal intrinsic, and internal relative FoRs. Avg represents the micro-average accuracy. CP refers to context with camera perspective, while RP denotes context with relatum perspective.
VISOR(%)
A-Split C-Split
Model cond (I) cond (R) cond (avg) cond (I) cond (R) cond (avg)
EI FoR ER FoR all
SD-1.5 $51.11$ $21.61$ $72.72$ $48.95$ $68.72$ $53.92$ $53.77$ $53.83$
SD-2.1 $57.97$ $21.49$ $79.46$ $54.10$ $75.39$ $\mathbf{60.06}$ $59.64$ $59.83$
Llama3-8B + GLIGEN $53.67$ $25.78$ $79.45$ $66.08$ $77.38$ $57.51$ $65.98$ $62.12$
Llama3-70B + GLIGEN $54.49$ $29.45$ $83.94$ $68.68$ $81.43$ $56.47$ $69.53$ $63.49$
Llama3-8B + SG + GLIGEN (Our) $57.46$ $27.96$ $85.42$ $\mathbf{71.14}$ $83.17$ $58.84$ $\mathbf{70.36}$ $\mathbf{65.15}$
Llama3-70B + SG + GLIGEN (Our) $56.54$ $30.59$ $\mathbf{87.13}$ $66.56$ $\mathbf{83.75}$ $56.77$ 70.0470.04{70.04}70.04 64.0664.06{64.06}64.06
Table 3: VISORcond score explained in Section 5.1 where $I$ refers to the Cow and Car cases where relatum has intrinsic directions, and $R$ refers to the Box and Pen cases where relatum lacks intrinsic directions, $avg$ is the micro-average of $I$ and $R$. EI and ER FoR represent the generated image considered corrected by EI or ER FoR.
5.1 Evaluation Metrics
QA. We report an accuracy (acc.) based on the correct answer defined as follows. Since the questions can have multiple correct answers, as explained in Section 3, the prediction is correct if it matches any valid answer. Additionally, we report the model’s bias distribution when FoR ambiguity exists. $I$is the percentage of correct answers when assuming an intrinsic FoR, while $R$is this percentage with a relative FoR assumption. Note that cases where both FoR assumptions lead to the same answer are excluded from the bias calculation.
T2I. We adopt spatialEval (Cho et al., 2023b) for evaluating T2I spatial ability. We modify it to account for FoR by converting relations to a camera perspective before passing them to spatialEval, which assumes this viewpoint. Accuracy is determined by comparing the bounding box and depth map of the relatum and locatum. For FoR ambiguity, a generated image is correct if it aligns with at least one valid FoR interpretation. We report results using VISORcond and VISORuncond (Gokhale et al., 2023) metrics. VISORcond evaluates spatial relations only when both objects appear correctly, aligning with our focus on spatial reasoning. While, VISORuncond evaluates the overall performance, including object creation errors.
5.2 Experimental Setting
QA. We use Llama3-70B (Llama, 2024), Qwen2-72B (Qwen Team, 2024), Qwen2VL-72B Wang et al. (2024), and GPT-4o (OpenAI, 2024) as backbone models. All models are evaluated with zero-shot, few-shot, CoT, and our SG+CoT prompting under temperature 0 to ensure reproducibility.
T2I. We select Stable Diffusion SD-1.5 and SD-2.1 (Rombach et al., 2021) as our stable diffusion models and GLIGEN(Li et al., 2023) as the layout-to-image backbone. For translating spatial descriptions into bounding box information, we use Llama3-8B and Llama3-70B, as detailed in Section 4.2. The same LLMs are used to generate spatial information for SG prompting. We generate four images to compute the VISOR score following Gokhale et al. (2023) Inference steps for all T2I models are set to 50. For the evaluation, we select grounding DINO (Liu et al., 2025) for object detection and DPT (Ranftl et al., 2021) for depth mapping, following VPEval Cho et al. (2023b). The experiments were conducted on two $A6000$ GPUs, totaling approximately $300$ GPU hours.
5.3 Results
RQ1. What is the LLM’s bias for the ambiguous FoR? Table 1 presents the QA results for the A-split. Ideally, a model extracting spatial relations without considering perspective should achieve 100% accuracy, as the context lacks a fixed perspective. However, our focus is not on this. We aim to assess LLMs’ bias by measuring how often they adopt a specific perspective when answering. In the Box and Pen cases, relatum properties do not introduce FoR ambiguity in directional relations, making the task pure extraction rather than reasoning. Thus, we focus on the $I$and $R$of the Cow and Car cases, which best reflect LLMs’ bias. Qwen2 achieves 80% accuracy across all experiments by selecting spatial relations directly from context, suggesting it may disregard the question’s perspective. This is supported by attention analysis in Appendix C. The attention maps show that most correct Qwen2 responses assume a shared perspective between context and question and pay low attention to the perspective mentioned in the question. GPT-4o exhibits similar bias in 0-shot and 4-shot settings but shifts toward intrinsic interpretation with CoT. This shift reduces accuracy in camera-perspective questions where FoR adaptation is more significant than relation extraction. Llama3-70B shows no strong preference, slightly favoring relative FoR, but this balanced assumption leads to lower performance due to increased reasoning demands. Qwen2VL exhibits a similar pattern but achieves higher accuracy, likely due to visual training. GPT-o4-mini with high visual reasoning shows the same pattern, which is later discussed as an additional experiment in Appendix J. Moreover, our experiments with humans, presumed to have prior visual knowledge, show strong reasoning ability, achieving over 90% accuracy even in scenarios where perspective preference requires FoR adaptation. In our results, humans slightly prefer intrinsic FoR over relative FoR. Their cultural background and recently seen examples may also influence their preference as discussed in Coventry et al. (2018). A more detailed analysis of human performance is provided in Appendix I. In summary, we conjecture that Qwen2 performs well by focusing on extraction without reasoning FoR, while other models attempt reasoning but often fail to reach correct conclusions, leading to lower accuracy.
Figure 3: Confusion matrices of predicted spatial relations by Llama3 and GPT-4o when adapting to FoR is required. 0-shot and SG-COT settings are reported here.
RQ2. Can LLMs adapt FoR when answering the questions? To address this RQ, we analyze QA results of C-split in Table 2, where context and question explicitly specify perspective. Results show that LLMs struggle with FoR conversion, achieving only up to 55.5% accuracy when the question uses the relatum’s perspective and the context uses the camera’s. We further examine how Llama3 and GPT-4o adapt FoR using the confusion matrix in Figure 3. Our findings reveal that pure-text LLM (Llama3) reverses left and right. This contrasts with humans when describing the spatial relation from a perspective in English, which reverses front and back while preserving lateral directions. This leads to poor adaptation to the camera perspective of Llama3. In contrast, large multimodal models like GPT-4o follow expected patterns, consistent with Zhang et al. 2025. While our GPT-4o results suggest some abilities to convert the relatum’s perspective into the camera’s with in-context learning (72% accuracy), the reverse transformation in the textual domain remains challenging (53% accuracy). A similar trend appears when comparing Qwen2 and Qwen2VL, as discussed in Appendix B. Qwen2VL performs better when shifting to the camera perspective but performs worse with reverse transformation, often failing to generate correct reasoning. This difficulty persists for several models when converting spatial relations from images to relatum’s perspective, as noted in Zhang et al. 2025. A possible reason is that the model often overlooks orientation tokens, critical for answering FoR questions, based on Attention analysis provided in Appendix C.
RQ3. How can an explicit FoR identification help spatial reasoning in QA? We compare CoT and CoT+SG results to assess the effect of FoR identification on LLMs’ spatial reasoning in QA. Based on C-Split results (Table 2), incorporating SG improves the model’s ability to identify the correct perspective from input expression ranging from 2.9% to 30% of cases where the context and question share the same perspective. These cases are easier as the models do not need FoR adaptation. Two notable exceptions emerge. First, Llama3 performs poorly on camera-perspective questions, and FoR identification via SG fails to improve its performance. This may be due to Llama3’s lack of visual training, which we speculate limits its FoR understanding. Second, Qwen2VL struggles with relatum-perspective reasoning, showing negative gains even with CoT. SG is less effective when context and question differ in perspective; while it helps identify the correct FoR in context, it does not enhance reasoning across perspectives. This limitation is evident in A-Split results (Table 1), where models only improve significantly when SG aligns their preference with the question’s perspective, as seen in Qwen2-72B and GPT-4o. SG identification results are reported in the Appendix H. Still, FoR identification improves overall spatial reasoning (see Avg. column for SG+CoT in Table 3).
RQ4. How can explicit FoR identification help spatial reasoning in visualization? We evaluate SG layout diffusion to assess the impact of incorporating FoR in image generation. We focus on VISORcond metric, which better reflects the model’s spatial understanding than the overall performance measured by VISORuncond, reported in Appendix F. Table 3 shows that adding spatial information and FoR classes (SG+GLIGEN) improves performance across all splits compared to the baseline models (GLIGEN). In particular, SG improved the model’s performance when expressions follow a relative FoR. This finding aligns with QA results in Table 1, where Llama3 prefers relative FoR in camera-perspective scenarios. In contrast, baseline diffusion models (SD-1.5 and SD-2.1) perform better for intrinsic FoR, even though GLIGEN is based on SD-2.1. This outcome may be due to GLIGEN’s reliance on bounding boxes for spatial configurations, which makes it struggle with intrinsic FoR due to the absence of object properties and orientation. Despite this bias, incorporating FoR information via SG-prompting improves performance across all FoR classes. We provide further analysis on SG for the layout generation in Appendix G.
6 Related Work
Frame of Reference in Cognitive Studies. The concept of the frame of reference in cognitive studies was introduced by Levinson 2003 and later expanded with more diverse spatial relations (Tenbrink, 2011). Subsequent research investigated the human preferences for specific FoR classes (Vukovic and Williams, 2015; Shusterman and Li, 2016; Ruotolo et al., 2016; Coventry et al., 2018). For instance, Ruotolo et al. 2016 examined how FoR influences scene memorization. Their study found that participants performed better when spatial relations were based on their position rather than external objects, highlighting a distinction between relative and intrinsic FoR.
Frame of Reference in AI. Several benchmarks have been developed to evaluate the spatial understanding of AI models in multiple modalities; for instance, textual QA (Shi et al., 2022; Mirzaee and Kordjamshidi, 2022; Rizvi et al., 2024), and text-to-image (T2I) benchmarks (Gokhale et al., 2023; Cho et al., 2023a, b). However, most of these benchmarks overlook FoR, assuming a single FoR for all instances despite its significance in cognitive studies. Recent vision-language studies have begun addressing this gap Liu et al. (2023); Zhang et al. (2025). For instance, Zhang et al. 2025 explores FoR ambiguity by evaluating spatial relations from camera-perspective images, with FoR specified in the question. In contrast, our work focuses on the spatial reasoning when dealing with multiple FoRs and when there are changes in perspective in explaining the context beyond the camera’s viewpoint. We further show that explicitly identifying FoR enhances spatial reasoning in QA and T2I.
7 Conclusion
Given the significance of spatial reasoning in AI, we introduce Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark to evaluate Frame of Reference comprehension in textual spatial expressions through question-answering and grounding in visual modality via diffusion models. Using this benchmark, we identify notable differences in FoR comprehension across LLMs and their struggle with questions that require adaptation between multiple FoRs. Moreover, the bias in FoR interpretations impacts the layout generation with LLMs for text-to-image models. To improve FoR comprehension, we propose Spatial-Guided prompting, which extracts spatial relations’ topological, distal, and directional type information in addition to FoR and includes this information in downstream task prompting. Employing SG improves performance in both QA tasks requiring FoR understanding and text-to-image generation.
8 Limitations
While we analyze LLMs’ shortcomings, our benchmark only highlights areas for improvement. The trustworthiness and reliability of the LLMs are still a research challenge. Our analysis is confined to the spatial reasoning domain and does not account for biases related to gender or race. However, we acknowledge that linguistic and cultural variations in spatial expression are not considered, as our study focuses solely on English. Extending this work to multiple languages could reveal important differences in FoR adaptation. Our analysis is still limited to the synthetic environment. Future research should consider the broader implications of the frame of reference of spatial reasoning in real-world applications. Additionally, our experiments require substantial GPU resources, limiting the selection of LLMs and constraining the feasibility of testing larger models. The computational demands also pose accessibility challenges for researchers with limited resources. We find no ethical concerns in our methodology or results, as our study does not involve human subjects or sensitive data.
Appendix A Additional details of FoREST Dataset
Figure 4: All 3d models used to generate visualizations for FoREST.
A.1 Dataset Statistics
The FoREST dataset statistic is provided in the Table 4 for A-split and Table 5 for C-split.
A.2 Details Creation of FoREST dataset
We define the nine categories of objects selected in our dataset as indicated below in Table 6. We select sets of locatum and relatum based on the properties of each class to cover four cases of frame of reference defined in Section 3.2. Notice that we also consider the appropriateness of the container; for example, the car should not contain the bus.
Based on the selected locatum and relatum. To create an A-split spatial expression, we substitute the actual locatum and relatum objects in the Spatial Relation template. After obtaining the A-split contexts, we create their counterparts using the perspective/topology clauses to make the counterparts in the C-split. Then, we obtain the I-A and I-C split by applying the directional template to the first occurrence of relatum when it has intrinsic directions. The directional templates are "that is facing towards," "that is facing backward," "that is facing to the left," and "that is facing to the right." All the templates are in the Table 7. We then construct the scene configuration from each modified spatial expression and send it to the simulator developed using Unity3D. Eventually, the simulator produces four visualization images for each scene configuration.
Category Object(s) Intrinsic Direction Possible Container
Small object without intrinsic directions umbrella, bag, suitcase, fire hydrant
Big object with intrinsic directions bench, chair
Big object without intrinsic direction water tank
Container box, container
Small animal chicken, dog, cat
Big animal deer, horse, cow, sheep
Small vehicle bicycle
Big vehicle bus, car
Tree tree
Table 6: All selected objects with two properties: intrinsic direction, affordance of being container
A.2.1 Simulation Details
The simulation starts with randomly placing the relatum into the scene with the orientation based on the given scene configuration. We randomly select the orientation by given scene configuration, [-40, 40] for front, [40, 140] for left, [140, 220] for back, and [220, 320] for right. Then, we create the locatum from the relatum position and move it in the spatial relation provided. If the frame of reference is relative, we move the locatum based on the camera’s orientation. Otherwise, we move it from the relatum’s orientation. Then, we check the camera’s visibility of both objects. If one of them is not visible, we repeat the process of generating the relatum until the correct placement is achieved. After getting the proper placement, we randomly choose the background from 6 backgrounds. Eventually, we repeat the procedures four times for one configuration.
A.2.2 Object Models and Background
For the object models and background, we find it from the unity assert store [^2] https://assetstore.unity.com. All of them are free and available for download. All of the 3D models used are shown in Figure 4.
{locatum} is in front of {relatum}
{locatum} is on the left of {relatum}
{locatum} is to the left of {relatum}
Spatial Relation Templates {locatum} is behind of {relatum}
{locatum} is back of {relatum}
{locatum} is on the right of {relatum}
{locatum} is to the right of {relatum}
within {relatum}
Topology Templates and inside {relatum}
and outside of {relatum}
from {relatum}’s view
relative to {relatum}
Perspective Templates from {relatum}’s perspective
from my perspective
from my point of view
relative to observer
{relatum} facing toward that camera
Orientation Templates {relatum}is facing away from the camera.
{relatum} facing left relative to the camera
{relatum} facing right relative to the camera
Table 7: All templates used to create FoREST dataset.
A.2.3 Textual templates
All manually created templates used to create FoREST are given in Table 7.
In the camera view, how is {locatum} positioned in relation to {relatum}?
Based on the camera perspective, where is the {locatum} from the {relatum}’s position?
Question Templates From the camera perspective, what is the relation of the{locatum} to the {relatum}?
Looking through the camera perspective, how does {locatum} appear to be oriented relative to {relatum}’s position?
Based on the camera angle, where is {locatum} located with respect to {relatum}’s location?
Table 8: All templates used to create FoREST dataset.
A.2.4 GPT template
We prompt GPT with an initial question using a specific template (the first template in Table 8) to generate four additional variations. Each generated template was manually reviewed and validated before being added to the set of questions in our corpus. All templates are listed in Table 8. Question templates for asking the spatial relation from relatum view are edited by replacing references from the camera to the relatum object.
Appendix B Comparison between Qwen2 and Qwen2-VL in QA setting
Figure 5: Confusion matrices of spatial relation answers when Qwen2 and Qwen2-VL must adapt FoR in the 0-shot and (SG+CoT) settings.
According to the A-split results in Table 1, Qwen2VL demonstrates a more balanced preference pattern and reasonable performance compared to Qwen2. This suggests that, unlike Qwen2, which often extracts spatial relations without accounting for perspective, Qwen2VL appears capable of reasoning about perspective shifts as well. This observation is further supported by the C-split results in Table 2, where Qwen2VL significantly outperforms Qwen2 in cases requiring adapting the spatial relations mentioned between different perspectives in most settings. These results support our hypothesis that visual training enhances the model’s ability to interpret perspective changes.
However, despite these improvements, Qwen2VL still struggles to provide reasoning from the relatum’s perspective on questions in CoT and SG+CoT settings. This suggests that while the model leverages visual training to reason about perspective and perform perspective shifts when necessary, its reasoning is more accurate when questions are framed from the camera’s viewpoint. However, it struggles with questions asked from the relatum’s perspective. This limitation is expected, as training data—particularly image captioning datasets—are typically annotated from a human, camera-based perspective, restricting the model’s ability to generalize across different viewpoints.
Appendix C Attention Analysis
Figure 6: Attention weights of input context for Question Answering in A-split. Considering the following semantic concepts: locatum mention in context, relation in context, relatum in context, relatum orientation, perspective in question, locatum in question, and relatum in question. Three patterns are considered: (1) incorrect answers, (2) correct answers assuming a shared perspective between context and question, and (3) correct answers assuming different perspectives.
C.1 Experiment Setting
We conducted an attention-based interpretability analysis using Qwen2 in a 0-shot setting, where only the context and question were provided. We focused on analyzing attention weights across key semantic elements in both A-split and C-split. We analyzed the attention weights for various groups of tokens that convey the following concepts: locatum mention occurring in the context, relation in context, relatum in context, perspective in context, relatum orientation, perspective in question, locatum in question, and relatum occurring in the question.
We categorized response patterns into three scenarios: (1) the model answers incorrectly, (2) the model answers correctly with the assumption that the same perspective holds for context and question, and (3) the model answers correctly with the assumption that context and question have different perspectives. Note that we separate the analysis for both splits because the conceptual tokens are different, that is, the perspective mentioned in the context is only for the C-split.
C.2 Results
A-split. According to Figure 6, we found that, for the correct responses, the model paid the least attention to the question’s perspective in the same perspective. Meanwhile, for the different perspective case, attention significantly increased to tokens related to the relation, the question perspective, and the relatum in the question. This result illustrates that these tokens played an important role in adapting spatial relations for perspective shifts. However, when the model increased attention weight on relatum orientation compared to other tokens, more errors occurred in answering the questions.
(a) Questions that require changing perspective.
(b) Questions that do not require changing perspective.
Figure 7: Attention weights of input context for Question Answering in C-split. Considering the following semantic concepts: locatum mention in context, relation in context, relatum in context, perspective in context, relatum orientation, perspective in question, locatum in question, and relatum in question. Two patterns are considered: (1) incorrect answers, (2) correct answers. We separate the questions that require and not-require perspective changing.
C-split. According to Figure 7, in cases that require a shift in perspective, we observed that the model pays significantly more attention to the spatial relation in context, the question perspective, and the relatum in the question concepts necessary for answering correctly. However, the model largely overlooks the orientation tokens, which also play a crucial role in resolving these questions. This lack of attention to orientation may majorly contribute to the model’s failures in perspective-shifting scenarios. In cases where the relation mentioned remains the same in both the context and the answer, we observed that the model, in incorrect predictions, tends to assign higher attention to the relatum in the question. In contrast, correct predictions are associated with more balanced attention across tokens. Maintaining moderate attention to all tokens appears to help the model consistently respond with the correct relation, repeating it from the context.
Appendix D Quantitative Examples
D.1 Example of QA Bias in the A-split
To illustrate example biases in the A-split, we present three types of model behavior. The first, shown in Figure 8, occurs when the model assumes a shared perspective between context and question, allowing it to extract spatial relations directly without FoR reasoning. This scenario is the most common behavior of Qwen2. The second, illustrated in Figure 9, arises when the model assumes the context always reflects the relatum’s perspective, favoring intrinsic over relative interpretation. Lastly, Figure 10 shows the model assuming the context always adopts the camera perspective.
Figure 8: The first example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model responds with the spatial relation in the context, assuming all questions and context have the same perspective.
Figure 9: The second example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model assumes the context has a relatum perspective.
Figure 10: The third example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model assumes the context has a camera perspective.
D.2 Incorrect QA reasoning in the C-split
Our results demonstrate that models struggle with converting between different perspectives when required to answer the questions, especially when the correct answer is either left or right. This difficulty contributes to their performance, as discussed in research question 2 in the paper (RQ2). To further illustrate LLMs’ struggle, we provide quantitative examples of failure cases, focusing specifically on the models’ most challenging instances in Figure 11.
Figure 11: An example of an incorrect SG+CoT case, where the model generates incorrect reasoning and mistakenly converts left to right.
Appendix E Template Variation in Question-Answering
We report average accuracy on the C-split across all question templates for Qwen2 and GPT-4o in Table 9. While results show notable variation across templates, no single template consistently outperforms others across all tasks. Template effectiveness depends on both the prompting technique and the underlying model. Simpler templates generally yield higher accuracy, though more complex expressions may reduce overall performance. Nevertheless, the comparative trend remains consistent, as SG+CoT often outperforms CoT in overall improvement across both models. Notably, SG prompting yields the smallest performance variation across templates.
Model T0 T1 T2 T3 T4 Avg. Highest $\Delta$
Qwen2 (0-shot) 70.23 61.36 69.35 65.55 67.97 66.91 8.87
Qwen2 (4-shot) 65.15 64.01 68.80 63.24 65.46 65.34 5.57
Qwen2 (CoT) 71.26 66.11 72.78 63.40 71.87 69.10 9.37
Qwen2 (SG + CoT) 71.15 70.92 72.89 70.11 71.69 71.36 2.78
GPT-4o (0-shot) 64.02 61.14 73.16 65.77 56.66 64.19 16.50
GPT-4o (4-shot) 65.85 52.23 70.55 61.64 58.33 61.78 18.31
GPT-4o (CoT) 80.98 69.65 84.81 83.24 74.98 78.78 15.16
GPT-4o (SG + CoT) 85.74 85.02 84.10 86.53 87.57 85.79 3.47
Table 9: Average accuracy across different templates. Ti refers to the generated QA template $i$ in Table 7. $\Delta$ measures the difference in accuracy between two different templates.
Appendix F VISOR-uncond Score
VISORuncond provides the overall spatial relation score, including images with object generation errors. Since it is less focused on evaluating spatial interpretation than VISORcond, which assesses explicitly the text-to-image model’s spatial reasoning, we report VISORuncond results here in the Table 10 rather than in the main paper. The results are similar to the pattern observed in VISORuncond that the based models(SD-1.5 and SD-2.1) perform better in the relative frame of reference, while the layout-to image models, i.e. GLIGEN, are better in the intrinsic frame of reference.
VISOR(%)
Model uncond (I) uncond (R) uncond (avg) uncond (I) uncond (R) uncond (avg)
I-A-Split I-C-Split
SD-1.5 $45.43$ $33.22$ $43.51$ $35.06$ $35.68$ $35.40$
SD-2.1 $\mathbf{62.87}$ $43.90$ $\mathbf{59.89}$ $\mathbf{45.98}$ $46.59$ $\mathbf{46.31}$
Llama3-8B + GLIGEN $46.74$ $38.16$ $45.39$ $33.98$ $39.36$ $36.89$
Llama3-70B + GLIGEN $54.33$ $46.89$ $53.17$ $38.04$ $46.04$ $42.37$
Llama3-8B + SG + GLIGEN (Our) $51.83$ $43.24$ $50.48$ $36.28$ $44.43$ $40.70$
Llama3-70B + SG + GLIGEN (Our) $58.92$ $\mathbf{47.44}$ $57.12$ $38.23$ $\mathbf{48.62}$ $43.86$
Table 10: VISORuncond score on the I-A-Split and I-C-Split where $I$ refer to the Cow Case and Car Case where relatum has intrinsic directions, and $R$ refer to the Box Case and Pen case where relatum lacks intrinsic directions, $avg$ is mirco-average of $I$ and $R$. cond and uncond are explained in Section 5.1.
Appendix G Analyze the improvements in SG-prompting for T2I.
Model Layout Layoutcond
Llama3-8B $85.26$ $88.84$
Llama3-8B + SG $85.04$ $88.86$
Llama3-70B $88.47$ $93.16$
Llama3-70B + SG $91.95$ $95.45$
Table 11: Layout accuracy where spatial relations are left or right relative to the camera. Layout is evaluated for all generated layouts in I-C split while Layoutcond uses the same testing examples as VISORcond.
To further explain improvements of SG-prompting in T2I task, we assess the generated bounding boxes in the I-C split for left and right relations relative to the camera since these can be evaluated using only bounding boxes without depth information. As seen is Table 11, our SG prompting improved Llama3-70B’s by $3.48\%$, while Llama3-8B saw a slight decrease of $0.22\%$. This evaluation was conducted on all generated layouts from the I-C split, which differs from the evaluation subset of images used for VISORcond in Table 3. We report the layoutcond score for a consistent evaluation in the same table. Layoutcond shows that Llama3-8B improves within the same evaluation subset with VISORcond. Overall, by incorporating FoR information through SG layout diffusion, Llama3 generates better spatial configurations, enhancing image generation performance.
Appendix H Frame of Reference Identification
We evaluate the LLMs’ performance in recognizing the FoR classes from given spatial expressions. The LLMs receive spatial expression, denoted as $T$, and output one FoR class, $FoR$, from the valid set of FoR classes, $For\in$external relative, external intrinsic, internal intrinsic, internal relative}. All in-context learning examples are in the Appendix K.
H.1 Experimental Setting
Zero-shot model. We follow the regular setting of zero-shot prompting. We only provide instruction to LLM with spatial context. The instruction prompt briefly explains each class of the FoR and candidate answers for the LLM. We called the LLM with the instruction prompt and $T$ to find $F$.
Few-shot model. We manually craft four spatial expressions for each FoR class. To avoid creating bias, each spatial expression is ensured to fit in only one FoR class. These expressions serve as examples of our few-shotsetting. We provide these examples in addition to the instruction as a part of the prompt, followed by $T$ and query $F$ from the LLM.
Chain-of-Thought (CoT) model. To create CoT (Wei et al., 2023) examples, we modify the prompt to require reasoning before answering. Then, we manually crafted reasoning explanations with the necessary information for each example used in few-shot. Finally, we call the LLMs, adding modified instructions to updated examples, followed by $T$ and query $F$.
Spatial-Guided Prompting (SG) model. We hypothesize that the general spatial relation types defined in Section 2 can provide meaningful information for recognizing FoR classes. For instance, a topological relation, such as “inside," is intuitively associated with an internal FoR. Therefore, we propose Spatial-Guided Prompting to direct the model in identifying the type of relations before querying $F$. We revise the prompting instruction to guide the model in considering these three aspects. Then, we manually explain these three aspects. We specify the relation’s origin from the context for direction relations, such as "the left direction is relative to the observer." We hypothesize that this information helps the model distinguish between intrinsic and relative FoR. Next, we specify whether the locatum is inside or outside the relatum for topological relations. This information should help distinguish between internal and external FoR classes. Lastly, we provide the potential quantitative distance, e.g., far. This quantitative distance further encourages identifying the correct topological and directional relations. Eventually, we insert these new explanations in examples and call the model with the updated instructions followed by $T$ to query $F$.
H.2 Evaluation Metrics
We report the accuracy of the model on the multi-class classification task. Note that the expressions in A-split can have multiple correct answers. Therefore, we consider the prediction correct when it is in one of the valid FoR classes for the given spatial expression.
Model inherently clear require template
CoT SG CoT SG
Llama3-70B 19.84 44.64 (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.80}$) 76.72 87.39 (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.67}$)
Qwen2-72B 58.20 84.22 (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 26.02}$) 88.36 93.86 (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.67}$)
GPT-4o 12.50 29.17 (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.67}$) 87.73 90.74 (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.01}$)
Table 12: The comparison between CoT and SG prompting in C-split separated by inherently clear / required template to be clear.
A-split C-Split
Model ER-C-Split EC-Split IC-Split IR-C-Split Avg.
Gemma2-9B (0-shot) $94.17$ $\mathbf{94.24}$ $35.98$ $53.91$ $57.66$ $60.45$
Gemma2-9B (4-shot) $59.58$ $55.89$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 38.34}$) $72.61$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 36.63}$) $74.22$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 20.31}$) $54.44$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.23}$) $64.29$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.84}$)
Gemma2-9B (CoT) $60.49$ $60.49$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 33.74}$) $60.54$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.57}$) $87.50$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 33.59}$) $54.03$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.63}$) $65.64$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 5.20}$)
Gemma2-9B (SG)(Our) $72.67$ $65.87$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 28.37}$) $65.54$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.57}$) $53.12$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.78}$) $\mathbf{95.97}$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 38.31}$) $70.13$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 9.68}$)
llama3-8B (0-shot) $60.21$ $32.20$ $90.11$ $75.78$ $0.00$ $49.52$
llama3-8B (4-shot) $60.14$ $47.77$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 15.58}$) $54.35$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 35.76}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22}$) $41.13$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 41.13}$) $60.81$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 11.29}$)
llama3-8B (CoT) $61.32$ $61.06$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 28.86}$) $97.28$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.17}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22}$) $36.29$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 36.29}$) $73.66$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.14}$)
llama3-8B (SG) (Our) $62.95$ $63.29$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 31.09}$) $94.57$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.46}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22}$) $43.55$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 43.55}$) $75.35$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 25.83}$)
llama3-70B (0-shot) $84.23$ $74.08$ $9.57$ $92.19$ $68.55$ $61.10$
llama3-70B (4-shot) $78.47$ $81.81$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.72}$) $64.89$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 55.33}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81}$) $75.81$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.26}$) $80.63$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 19.53}$)
llama3-70B (CoT) $69.11$ $72.05$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 2.03}$) $97.07$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 87.50}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81}$) $79.44$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.89}$) $87.14$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 26.04}$)
llama3-70B (SG) (Our) $76.50$ $78.21$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.12}$) $97.61$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 88.04}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81}$) $72.18$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.63}$) $87.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 25.90}$)
Qwen2-7B (0-shot) $83.64$ $79.97$ $59.24$ $77.34$ $40.73$ $64.32$
Qwen2-7B (4-shot) $61.12$ $50.52$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 29.45}$) $65.76$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 6.52}$) $93.75$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.41}$) $56.05$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 15.32}$) $66.52$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 2.20}$)
Qwen2-7B (CoT) $72.12$ $70.81$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 9.16}$) $63.80$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.57}$) $99.22$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 21.88}$) $51.61$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.89}$) $71.36$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.04}$)
Qwen2-7B (SG) $70.61$ $68.00$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 11.98}$) $71.20$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 11.96}$) $88.28$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.94}$) $57.26$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.53}$) $71.18$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 6.86}$)
Qwen2-72B (0-shot) $64.46$ $62.70$ $100.00$ $100.00$ $39.11$ $75.45$
Qwen2-72B (4-shot) $79.12$ $78.73$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.03}$) $99.35$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.65}$) $87.50$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 12.50}$) $87.10$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 47.98}$) $88.17$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 12.72}$)
Qwen2-72B (CoT) $88.54$ $88.87$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 26.18}$) $89.57$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 10.43}$) $93.75$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 6.25}$) $83.47$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 44.35}$) $88.91$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 13.46}$)
Qwen2-72B (SG) $90.51$ $90.18$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 27.49}$) $93.26$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 6.74}$) $98.44$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 1.56}$) $85.08$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 45.97}$) $91.74$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.29}$)
GPT3.5 (0-shot) $83.11$ $88.15$ $17.50$ $70.31$ $41.13$ $54.27$
GPT3.5 (4-shot) $61.25$ $48.95$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 39.20}$) $62.72$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 45.22}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.69}$) $28.63$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 12.50}$) $60.07$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 5.80}$)
GPT3.5 (CoT) $66.55$ $66.62$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 21.53}$) $96.85$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 79.35}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.69}$) $50.81$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 9.68}$) $78.57$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.30}$)
GPT3.5 (SG) (Our) $70.61$ $73.30$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 14.86}$) $92.93$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 75.43}$) $99.22$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 28.91}$) $49.19$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 8.06}$) $78.66$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.39}$)
GPT4o (0-shot) $73.82$ $71.27$ $98.80$ $100.00$ $70.56$ $85.16$
GPT4o (4-shot) $66.23$ $67.87$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.40}$) $98.70$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.11}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00}$) $78.63$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 8.06}$) $86.30$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.14}$)
GPT4o (CoT) $72.44$ $72.77$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.51}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.20}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00}$) $73.79$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.23}$) $86.64$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.48}$)
GPT4o (SG) (Our) $76.44$ $74.67$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.40}$) $97.72$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 1.09}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00}$) $68.55$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 2.02}$) $85.23$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.08}$)
Table 13: Accuracy results report from FoR Identification with LLMs. The correct prediction is one of the valid FoR classes for the given spatial expression. All FoR classes are external relative (ER), external intrinsic (EI), internal intrinsic (II), and internal relative (IR).
(a) Results of Cow Case in A-Split.
(b) Results of Car Case in A-Split.
(c) Results of Box Case in A-Split.
(d) Results of Pen Case in A-Split.
Figure 12: Red shows the wrong FoR identifications, and green shows the correct ones. The dark color is for relative FoRs, while the light color is for intrinsic FoRs. The round shape is for the external FoRs, while the square is for internal FoRs. The depth of the plots shows the four FoRs, i.e., external relative, external intrinsic, internal intrinsic, and internal relative, from front to back.
(a) Results of Cow Case in A-Split.
(b) Results of Car Case in A-Split.
(c) Results of Box Case in A-Split.
(d) Results of Pen Case in A-Split.
Figure 13: Red shows the wrong FoR identifications, and green shows the correct ones. The dark color is for relative FoRs, while the light color is for intrinsic FoRs. The round shape is for the external FoRs, while the square is for internal FoRs. The depth of the plots shows the four FoRs, i.e., external relative, external intrinsic, internal intrinsic, and internal relative, from front to back. This plot is the result of the rest of LLMs.
H.3 Results
H.3.1 FoR Inherently Bias in LLMs
C-spilt. The zero-shot setting reflects the LLMs’ inherent bias in identifying FoR. Table 13 presents the accuracy for each FoR class in C-split, where sentences explicitly include information about topology and perspectives. We found that some models strongly prefer specific FoR classes. Notably, Gemme2-9B achieves a near-perfect accuracy on external relative FoR but performs poorly on other classes, especially external intrinsic, indicating a notable bias towards external relative. In contrast, GPT4o and Qwen2-72B perform exceptionally in both intrinsic FoR classes. However, they perform poorly in the relative FoRs. We also
A-spilt. We examine the FoR bias in the A-split. Based on the results in Table 13, we plotted the top-3 models’ results (Gemma2-9B, Llama3-70B, and GPT4o) for a more precise analysis in Figures 12. The plots show the frequencies of each FoR category. According to the plot, Gemma and GPT have strong biases toward external relative and external intrinsic, respectively. This bias helps Gemma2 perform well in the A-split since all spatial expressions can be interpreted as external relative. However, GPT4o’s bias leads to errors when intrinsic FoRs aren’t valid, as in the Box and Pen cases (see plots (c) and (d)). Llama3 exhibits different behavior, showing a bias based on the relatum’s properties, specifically the relatum’s affordance as a container. In cases where relatum cannot serve as containers, i.e., Cow and Pen cases, Llama3 favors external relative. Conversely, Llama3 tends to favor external intrinsic when the relatum has the potential to be a container.
H.3.2 Behavior with ICL variations
C-spilt. We evaluate the models’ behavior under various in-context learning (ICL) methods. As observed in Table 13, the few-shot method improves the performance of the zero-shot method across multiple LLMs by reducing their original bias toward specific classes. Reducing the bias, however, lowers the performance in some cases, such as the performance of Gemma 2 in ER class. One noteworthy observation is that while the CoT prompting generally improves performance in larger LLMs, it is counterproductive in smaller models for some FoR classes. This suggests that the smaller models have difficulty inferring FoR from the longer context. This negative effect also appears in SG prompting, which uses longer explanations. Despite performance degradation in particular classes of small models, SG prompting performs exceptionally well across various models and achieves outstanding performance with Qwen2-72B. We further investigate the performance of CoT and SG prompting. As shown in Table 12, CoT exhibits a substantial difference in performance between contexts with inherently clear FoR and contexts requiring the template to clarify FoR ambiguity. This implies that CoT heavily relies on the specific template to identify FoR classes. In contrast, SG prompting demonstrates a smaller gap between these two scenarios and significantly enhances performance over CoT in inherently clear FoR contexts. Therefore, guiding the model to provide characteristics regarding topological, distance, and directional types of relations improves FoR comprehension.
A-spilt. We use the same Figure 12 to observe the behavior when applying ICL. The A-split shows minimal improvement with ICL variations, though some notable changes are observed. With few-shot, all models show a strong bias toward external intrinsic FoR, even when the relatum lacks intrinsic directions, i.e., Box and Pen cases. This bias appears even in Gemma2-9B, which usually behaves differently. This suggests that the models pick up biases from the examples despite efforts to avoid such patterns. However, CoT reduces some bias, leading LLMs to revisit relative, which is generally valid across scenarios. In Gemma2, the model predicts relative FoR where the relatum has intrinsic directions, i.e., Cow and Car cases. Llama3 behaves similarly in cases where the relatum cannot act as a container, i.e., Cow and Pen cases. GPT4o, however, does not depend on the relatum’s properties and shows slight improvements across all cases. Unlike CoT, our SG prompting is effective in all scenarios. It significantly reduces biases while following a similar pattern to CoT. Specifically, SG prompting increases external relative predictions for Car and Cow in Gemma2-9B, and for Cow and Pen in Llama3-70B. Nevertheless, GPT4o shows only a slight bias reduction. However, Our proposed method improves the overall performance of most models, as shown in Table 13. The Llama3-70B behaviors are also seen in LLama3-8B and GPT3.5. The plots for LLama3-8B and GPT3.5 are in Figure 13.
H.3.3 Experiment with different temperatures
We conducted additional experiments to further investigate the impact of temperature on the biased interpretation of the model in the A-split of our dataset. As presented in Table 14, comparing distinct temperatures (0 and 1) revealed a shift in the distribution. The frequencies of the classes experienced a change of up to 10%. However, the magnitude of this change is relatively minor, and the relative preferences for most categories remained unchanged. Specifically, the model exhibited the highest frequency responses for the cow, car, and pen cases, even with higher frequencies in certain settings. Consequently, a high temperature does not substantially alter the diversity of LLMs’ responses to this task, which is an intriguing finding.
Model ER EI II IR
temp-0 temp-1 temp-0 temp-1 temp-0 temp-1 temp-0 temp-1
Cow Case
0-shot 75.38 87.12 23.86 12.50 0.76 0.13 0.00 0.25
4-shot 0.00 15.66 100.00 84.34 0.00 0.00 0.00 0.00
CoT 31.82 49.87 68.18 49.87 0.00 0.13 0.00 0.13
SG 51.39 70.45 48.61 29.42 0.00 0.00 0.00 0.13
Box Case
0-shot 22.50 41.67 77.50 58.33 0.00 0.13 0.00 0.25
4-shot 0.00 0.00 100.00 100.00 0.00 0.00 0.00 0.00
CoT 0.00 5.83 100.00 94.17 0.00 0.00 0.00 0.00
SG 11.67 33.33 88.33 66.67 0.00 0.00 0.00 0.00
Car Case
0-shot 55.20 68.24 49.01 31.15 0.79 0.61 0.00 0.00
4-shot 0.60 5.94 99.40 94.06 0.00 0.00 0.00 0.00
CoT 19.64 38.52 80.16 61.27 0.20 0.20 0.00 0.00
SG 44.25 56.97 55.75 43.03 0.00 0.00 0.00 0.00
Pen Case
0-shot 90.62 96.88 9.38 3.12 0.00 0.61 0.00 0.00
4-shot 0.00 7.03 100.00 92.97 0.00 0.00 0.00 0.00
CoT 17.19 28.91 82.81 71.09 0.20 0.20 0.00 0.00
SG 48.31 57.81 54.69 42.19 0.00 0.00 0.00 0.00
Table 14: The results between two different temperatures of Llam3-70B on the A-spilt of FoREST. The number shows the percentage frequency of responses from the model.
Appendix I Study of Human bias in A-spilt
I.1 Experiment Setting
We provide a Google Form with instructions in Listing 1, followed by 150 randomly ordered context-question pairs. These are sampled based on relatum category: 25 questions per case from the A-split, covering four cases (Cow, Car, Box, Pen) for camera-perspective questions, and two cases (Cow, Car) for relatum-perspective questions, as Box and Pen lack intrinsic direction. No time constraints were imposed, and all participants were compensated at the standard research assistant rate. All participants were informed about the study details at the end and asked for consent to use their responses in the analysis. After acquiring the results, we utilize the same metric defined in Section 5.1. Then, we average the accuracy of each participant to derive the human baseline based on this study.
I.2 Results
Individual human results are shown in Table 15. These results indicate that humans rely heavily on their assumptions when interpreting ambiguous frame-of-reference (FoR) contexts. Most participants favored the intrinsic FoR in A-split scenarios, though this preference varied across individuals, suggesting that background or prior visual experience may influence FoR interpretation. This aligns with findings from cognitive studies on how humans describe spatial relations that recently seen examples may influence Coventry et al. (2018). However, our results contrast with other studies suggesting that humans tend to adopt an egocentric (relative) FoR in scenes with fewer landmarks and an allocentric (intrinsic) FoR in those with more landmarks Forster et al. (2023). Since our study does not provide visual input, we cannot infer how participants mentally reconstruct scenes to answer the questions with only two objects in the scene. The next notable point is that while Box and Pen cases appear straightforward for models (Table 1), humans may still struggle with questions in these categories. Humans do not exhibit the same limitations as LLMs in perspective-shifting tasks. When the context and question differ in perspective, humans still achieve over 90% accuracy, whereas LLMs often struggle with this type of reasoning.
Instruction:
You will be provided with a scene description that describes the spatial relationship between two objects.
The scene description may include object orientation, such as A is facing toward the camera, to indicate the direction A is facing.
You will then receive a question asking about the spatial relationship between A and B from either the camera’s perspective or an object’s perspective.
Your task is to answer the question based on your understanding of the given spatial relationship.
Listing 1: Insturction for collecting human results on QA using A-split of FoREST dataset.
Camera perspective Relatum perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
RIAcc. RIAcc. Acc. Acc. Acc. RIAcc. RIAcc. Acc.
P #1 $22.7$ $\mathbf{77.3}$ $88.0$ $8.3$ $\mathbf{91.7}$ $96.0$ $92.0$ $96.0$ $93.0$ $20.8$ $\mathbf{79.2}$ $96.0$ $4.2$ $\mathbf{95.8}$ $96.0$ $96.0$
P #2 $8.7$ $\mathbf{91.3}$ $92.0$ $0.0$ $\mathbf{100.0}$ $96.0$ $36.0$ $56.0$ $70.0$ $20.0$ $\mathbf{80.0}$ $100.0$ $16.7$ $\mathbf{83.3}$ $96.0$ $98.0$
P #3 $\mathbf{78.3}$ $21.7$ $92.0$ $\mathbf{75.0}$ $25.0$ $96.0$ $88.0$ $96.0$ $93.0$ $\mathbf{83.3}$ $16.7$ $96.0$ $\mathbf{87.5}$ $12.5$ $96.0$ $96.0$
Human-baseline $36.6$ $\mathbf{63.4}$ $90.7$ $27.8$ $\mathbf{72.2}$ $96.0$ $72.0$ $82.7$ $85.3$ $41.4$ $\mathbf{58.6}$ $97.3$ $36.1$ $\mathbf{63.9}$ $96.0$ $96.7$
Table 15: QA accuracy in the A-Split of human study. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Appendix J Additional Results with GPT-o4-mini-high
We include results from GPT-4o-mini-high, a model configured for enhanced reasoning, for comparison with our previous results in Table 1 and Table 2. In this experiment, the reasoning-effort parameter is set to high. As shown in Table 16, the model exhibits a balanced preference between relative and intrinsic interpretations. Similar to Qwen2VL, it demonstrates strong reasoning ability, achieving high accuracy even when FoR adaptation is required due to its visual reasoning capability. Notably, GPT-o4-mini-high performs well on questions from the relatum perspective, even outperforms the GPT-4o (SG+CoT) in Table 17. However, its performance declines on camera-perspective questions. We believe this is possibly due to training emphasizing alternative perspectives over the camera view, which may reduce their performance from the camera perspective. Therefore, even with a strong reasoning model, the perspective shift remains challenging in the textual domain.
Camera perspective Relatum perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
RIAcc. RIAcc. Acc. Acc. Acc. RIAcc. RIAcc. Acc.
GPT-4o (1) $\mathbf{84.3}$ $15.3$ $94.5$ $\mathbf{88.5}$ $11.0$ $97.3$ $99.2$ $99.8$ $95.6$ $21.6$ $\mathbf{78.0}$ $91.6$ $16.1$ $\mathbf{83.5}$ $90.5$ $91.4$
GPT-4o (2) $\mathbf{69.0}$ $30.6$ $76.6$ $\mathbf{80.3}$ $19.2$ $89.5$ $100.0$ $100.0$ $81.5$ $29.0$ $\mathbf{70.5}$ $74.7$ $30.9$ $\mathbf{68.7}$ $77.5$ $75.1$
GPT-4o (3) $41.5$ $\mathbf{58.3}$ $92.3$ $38.2$ $\mathbf{61.6}$ $91.0$ $100.0$ $99.8$ $93.2$ $33.9$ $\mathbf{65.8}$ $93.9$ $32.0$ $\mathbf{67.6}$ $93.9$ $93.9$
GPT-4o (4) $26.0$ $\mathbf{73.9}$ $79.2$ $27.7$ $\mathbf{72.1}$ $79.4$ $96.7$ $94.3$ $81.4$ $16.2$ $\mathbf{83.4}$ $95.5$ $19.2$ $\mathbf{80.4}$ $94.8$ $95.4$
o4-mini-high $\mathbf{68.0}$ $31.6$ $92.0$ $\mathbf{69.9}$ $29.7$ $94.7$ $100.0$ $99.4$ $93.4$ $\mathbf{58.5}$ $41.3$ $86.9$ $\mathbf{50.2}$ $49.5$ $93.3$ $87.8$
Human-baseline $36.6$ $\mathbf{63.4}$ $90.7$ $27.8$ $\mathbf{72.2}$ $96.0$ $72.0$ $82.7$ $85.3$ $41.4$ $\mathbf{58.6}$ $97.3$ $36.1$ $\mathbf{63.9}$ $96.0$ $96.7$
Table 16: Additional results of QA accuracy in the A-Split with GPT-o4-mini-high. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Camera perspective Relatum perspective
Model ER (CP) EI (RP) II (RP) IR (CP) Avg. ER (CP) EI (RP) II (RP) IR (CP) Avg.
GPT-4o (0-shot) $79.7$ $45.1$ $39.5$ $90.2$ $64.2$ $46.9$ $88.5$ $98.2$ $34.8$ $67.5$
GPT-4o (4-shot) $68.0$ $52.6$ $60.7$ $74.1$ $61.8$ $44.9$ $\mathbf{98.2}$ $\mathbf{100.0}$ $37.5$ $71.2$
GPT-4o (CoT) $81.7$ $\mathbf{76.1}$ $\mathbf{82.4}$ $71.5$ $78.8$ $53.0$ $91.1$ $90.6$ $50.8$ $71.9$
GPT-4o (SG+CoT) $\mathbf{97.9}$ $72.2$ $72.7$ $\mathbf{93.4}$ $\mathbf{85.8}$ $48.9$ $96.3$ $95.9$ $36.1$ $71.8$
o4-mini-high $68.7$ $73.1$ $72.1$ $79.0$ $71.4$ $\mathbf{78.9}$ $91.9$ $93.9$ $\mathbf{62.7}$ $\mathbf{84.5}$
Table 17: Additional results of QA accuracy in the C-Split with GPT-o4-mini-high. ER, EI, II, and IR denote external relative, external intrinsic, internal intrinsic, and internal relative FoRs. Avg represents the micro-average accuracy. CP refers to context with camera perspective, while RP denotes context with relatum perspective.
Appendix K LLM Prompt
K.1 FoR Identification
We provide the prompting for each in-context learning. The prompting for zero-shot and few-shot is provided in Listing 2. The instruction answer for these two in-context learning is “Answer only the category without any explanation. The answer should be in the form of {Answer: Category.}"
For the Chain of Thought (CoT), we only modified the instruction answer to “Answer only the category with an explanation. The answer should be in the form of {Explanation: Explanation Answer: Category.}" Similarly to CoT, we only modified the instruction answer to “Answer only the category with an explanation regarding topological, distance, and direction aspects. The answer should be in the form of {Explanation: Explanation Answer: Category.}", respectively. The example responses are provided in Listing 5 for Spatial Guided prompting.
# Instruction to find frame of reference class of given context
Instruction:
You specialize in language and spatial relations, specifically in the frame of context (multiple perspectives in the spatial relation). Identify the frame of reference category given the following context. There are four classes of the frame of reference (external intrinsic, internal intrinsic, external relative, internal relative). Note that the intrinsic direction refers to whether the model has the front/back by itself. (Example: a bird, human. Counter Example: a ball, a box). "
External intrinsic. The spatial description of an entity A relative to another entity B, where (1) A is not contained by B, (2) the spatial relation is based on B’s facing orientation (intrinsic direction) if B has one.
Internal intrinsic. The spatial description of an entity A relative to another entity B, where (1) A is contained by B, (2) the spatial relation is based on B’s facing orientation (intrinsic direction) if B has one.
External relative. The spatial description of an entity A relative to another entity B, where (1) A is not contained by B, (2) The spatial relation is relative to entity C, where C is not involved in the spatial relation and only provides the perspective. An example of C is the camera.
Internal relative. The spatial description of an entity A relative to another entity B, where (1) A is contained by B, (2) The spatial relation is relative to entity C, where C is not involved in the spatial relation and only provides the perspective. An example of C is the camera.
{Instruction answer}
# Normal Instruction answer: Find the frame of reference category without any explanations. The answer should be in the form of Answer: Category.
# COT Instruction answer: Find the frame of reference category with an explanation. The answer should be in the form of Explanation: Explanation. Answer: Category.
# SG Instruction answer: Find the frame of reference category with an explanation. Additionally, include three more explanations relating to topology, distance, and direction from the given context. The answer should be in the form of \"Explanation: Explanation. Topology: Explanation. Distance: Explanation. Direction: Explanation. Answer: Category.
Context: {spatial exprssion}
Listing 2: Prompt for finding the frame of reference class of given context.
## These are four examples used in few-shot, CoT, and SG
# Example 1: The bird is outside the car and in front of the car relative to the car.
# Few-shot response: Answer: external intrinsic
# CoT response: Explanation: The car does not contain the bird, and the spatial relation is relative to the car’s facing orientation. The intrinsic direction of the car is relevant in this context, as the description \"in front of\" implies a direction relative to the sheep’s front or back. Answer: external intrinsic.
# SG response: Explanation: The car does not contain the bird, and the spatial relation is relative to the car’s facing orientation. The intrinsic direction of the car is relevant in this context, as the description \"in front of\" implies a direction relative to the sheep’s front or back. Topology: The bird is outside the car. Distance: There is little distance between the bird and the car. Direction: The bird is in the front direction of the car (entity B). Answer: external intrinsic.
# Example 2: The bird is inside of the car and left of the car from the car’s perspective.
# Few-shot response: Answer: internal intrinsic
# CoT response: Explanation: In this scenario, the bird is contained within the car, meaning it is an internal relationship. Additionally, the description specifies the bird’s position relative to the car, indicating that the spatial relation is relative to the car itself, considering its intrinsic direction. Answer: internal intrinsic.
# SG response: Explanation: In this scenario, the bird is contained within the car, meaning it is an internal relationship. Additionally, the description specifies the bird’s position relative to the car, indicating that the spatial relation is relative to the car itself, considering its intrinsic direction. Topological: The bird is inside the car. Distance: The bird is in the same location as the car. Direction: The bird is in the left direction of the car (entity B). Answer: internal intrinsic.
# Example 3: The box is inside and at the back of the room.
# Few-shot response: Answer: internal relative
# CoT response: Explanation: In this context, the box (entity A) is described as inside the room (entity B), fulfilling the first condition for an internal frame of reference. However, the spatial relation given (\" at the back\") cannot originate from the room since the room does not have intrinsic direction, implying it is relative to an external observer or reference point (C), such as a camera or a person viewing the scene. Answer: internal relative
# SG response: Explanation: In this context, the box (entity A) is described as inside the room (entity B), fulfilling the first condition for an internal frame of reference. However, the spatial relation given (\" at the back\") cannot originate from the room since the room does not have intrinsic direction, implying it is relative to an external observer or reference point (C), such as a camera or a person viewing the scene. Topological: The box is inside the room. Distance: The box is in the same location as the room. Direction: From the observer’s perspective, the box is located at the back of the room. Answer: internal relative
# Example 4: A phone is on the left of a tablet from my perspective.
# Few-shot response: Answer: external relative
# CoT response: Explanation: In this context, the spatial relation is described from the perspective of an observer (C) who is not involved in the spatial relation. The phone (A) is not contained by the tablet (B), and the spatial relation \"left of\" is relative to the observer’s perspective, not the tablet’s facing orientation. Answer: external relative.
# SG response: Explanation: In this context, the spatial relation is described from the perspective of an observer (C) who is not involved in the spatial relation. The phone (A) is not contained by the tablet (B), and the spatial relation \"left of\" is relative to the observer’s perspective, not the tablet’s facing orientation. Topological: The phone is not contained by the tablet. Distance: There is some distance between the phone and the tablet. Direction: From the observer’s perspective, the phone is located to the left of the tablet. Answer: external relative.
Listing 3: Spatial expression examples using for few-shots, Chain-of-Thought (CoT), and Spatial Guide (SG) prompting
K.2 Question Answering Example
#Context: The bird is outside the car and in front of the car relative to the car. The car is facing toward the camera.
#Question: Based on the camera’s perspective, where is the bird from the car position in the scene?
#SG information (used in SG + CoT): Frame of reference Explanation: The car does not contain the bird, and the spatial relation is relative to the car’s facing orientation. The intrinsic direction of the car is relevant in this context, as the description \"in front of\" implies a direction relative to the sheep’s front or back. Topology: The bird is outside the car. Distance: There is little distance between the bird and the car. Direction: The bird is in the front direction of the car (entity B). Frame of reference: external intrinsic.
#Normal response: Answer: front
#CoT / (SG + COT) response:Explanation: Based on the context, the bird’s position is in the front direction of the car. The car is facing toward the camera. Then, the car’s front direction is the camera’s front direction. Therefore, the bird’s position is in front of the car’s position from the camera’s perspective. Answer: front
#Context: The bird is inside the car and left of the car from the car’s perspective. The car is facing to the right relative to the camera. #Question: Based on the camera’s perspective, where is the bird from the car’s position?
#SG information (used in SG + CoT): Frame of reference Explanation: In this scenario, the bird is contained within the car, meaning it is an internal relationship. Additionally, the description specifies the bird’s position relative to the car, indicating that the spatial relation is relative to the car itself, considering its intrinsic direction. Topological: The bird is inside the car. Distance: The bird is in the same location as the car. Direction: The bird is in the left direction of the car (entity B). Frame of reference: internal intrinsic.
#Normal response: Answer: back
#CoT / (SG + COT) response: Explanation: Based on the context, the bird’s position is in the left direction of the car. The car is facing to the right. Then, the car’s left direction is the camera’s back direction. Therefore, the bird’s position is to the back of the car’s position from the camera’s perspective. Answer: back
Context: The box is inside and to the right of the room from the observer’s perspective.
Question: From the observer’s perspective, what is the spatial relation of the box to the room?
#SG information (used in SG + CoT): Frame of reference Explanation: In this context, the box (entity A) is described as inside the room (entity B), fulfilling the first condition for an internal frame of reference. However, the spatial relation given (\" at the back\") cannot originate from the room since the room does not have intrinsic direction, implying it is relative to an external observer or reference point (C), such as a camera or a person viewing the scene. Topological: The box is inside the room. Distance: The box is in the same location as the room. Direction: From the observer’s perspective, the box is located at the back of the room. Frame of reference: internal relative.
#Normal response: Answer: right
#CoT / (SG + COT) response: Explanation: Based on the context, the box is to the right of the room from the camera’s direction. Therefore, the box’s position is to the right of the room’s position from the observer’s perspective. Answer: right
Context: A phone is to the left of a tablet from my perspective. The tablet is facing to the right. Question: From my perspective, what is the spatial relation of the phone to the tablet?
#SG information (used in SG + CoT): Frame of Reference Explanation: In this context, the spatial relation is described from the perspective of an observer (C) who is not involved in the spatial relation. The phone (A) is not contained by the tablet (B), and the spatial relation \"left of\" is relative to the observer’s perspective, not the tablet’s facing orientation. Topological: The phone is not contained by the tablet. Distance: There is some distance between the phone and the tablet. Direction: From the observer’s perspective, the phone is located to the left of the tablet. Frame of Reference: external relative.
#Normal response: Answer: left
#CoT / (SG + COT) response: Explanation: Based on the context, the phone is to the left of the tablet from my perspective. The direction of the tablet is not relevant in this context since the left relation is from my perspective. Therefore, from my perspective, the phone is to the left of the tablet. Answer: left
Listing 4: Spatial expression examples using for few-shots, Chain-of-Thought (CoT), and Spatial Guide (SG) prompting for question-answering.
K.3 Text to Layout
# Instruction for generating bounding box
Your task is to generate the bounding boxes of objects mentioned in the caption.
The image is size 512x512. The bounding box should be in the format of (x, y, width, height). Please considering the frame of reference of caption and direction of reference object if possible. If needed, you can make the reasonable guess.
Listing 5: Prompt for generating bounding coordinates to use as the layout for layout-to-image models.